OpenStudy (anonymous):

where does cot x equal 1 on a the graph y=cotx on the interval [0, 2π]

OpenStudy (anonymous):

cotan(x)=1 is the same as tan(x)=1

OpenStudy (anonymous):

cotan x = (1/tan) x If cotan x = 1, then 1/1 = cotan x (1/1)/1 = 1

OpenStudy (anonymous):

You should be able to use a unit circle, or use arctan

OpenStudy (anonymous):

arctan(1) = 45 degrees or 225 degrees